OpenStudy (loser66):

Use induction to prove \(\sum_{k=1}^n \dfrac{1}{k^2}\leq 2-\dfrac{1}{n}\) for all positive integers n. Please, help

OpenStudy (loser66):

HS: \(\sum_{k=1}^n \dfrac{1}{k^2}\leq 2 - \dfrac{1}{n}\) Induction step \(\sum_{k=1}^{n+1} \dfrac{1}{k^2}= \sum_{k=1}^n \dfrac{1}{k^2}+\dfrac{1}{(n+1)^2}\leq 2-\dfrac{1}{(n+1)^2}\)

OpenStudy (loser66):

oh, the last one is not ^2 \(\sum_{k=1}^{n+1} \dfrac{1}{k^2}= \sum_{k=1}^n \dfrac{1}{k^2}+\dfrac{1}{(n+1)^2}\leq 2-\dfrac{1}{(n+1)}\)

OpenStudy (loser66):

This is my attempt: the first part of the sum is \(\leq 2 - (1/n)\) , hence the LHS is \(\leq 2 -(1/n) + (1/n+1)^2\)

OpenStudy (loser66):

So, we need prove \(2-\dfrac{1}{n} +\dfrac{1}{(n+1)^2}\leq 2- \dfrac{1}{n+1}\)

OpenStudy (loser66):

Then I stuck

ganeshie8 (ganeshie8):

clear the denominators and simplify

ganeshie8 (ganeshie8):

it simplifies nicely and you end up with \(1\ge 0\)

OpenStudy (loser66):

Yes, I did, but those are denominators, hence when I take reciprocal, the signs switch around.

OpenStudy (loser66):

oh, let me try again. :)

ganeshie8 (ganeshie8):

\(2-\dfrac{1}{n} +\dfrac{1}{(n+1)^2}\leq 2- \dfrac{1}{n+1}\) \(\iff\) \(\dfrac{1}{n} -\dfrac{1}{(n+1)^2}\ge \dfrac{1}{n+1}\) \(\iff\) \((n+1)^2-n\ge n(n+1)\)

OpenStudy (loser66):

Thanks a lot.

OpenStudy (loser66):

I got it. Thank you so much

ganeshie8 (ganeshie8):

np